MARS DESERT RESEARCH STATION

Sol Summary – April 18th

MDRS Daily Summary Report for sol 8

Summary Title: 9th EVA

Mission Status:  Good

Sol Activity Summary:

Hab
8.00:   Breakfast
10.00: Muscle wasting exercises
13.00: Lunch

GreenHab
/

ScienceDome
Experiment reports redaction

EVA
10.30: UTM NAD27 12S 4251500-518500
Telecommunication relays test
Diffractometer study: data collection

Weather:
Sunny and strong wind

Crew Physical Status
Everyone is fine

Look Ahead Plan:
EVA + Experiments

Reports to be file: Operation Report, Daily Summary Report, Commander Report, EVA 9 Report

Support Requested:

Astronomy Report – April 17th

Hello Peter,

We tried yesterday night to focus on shining stars. We used the black focus first, then the gold one, we understand the way it is supposed to work, but here is the problem : the image is never sharp, and when we defocus the star become « twisted » in one direction. To illustrate the problem, I joined a sketch of our problem. The quality of the draw is awfull, I’m sorry, we can’t use internet so I did what I could. Mathieu thinks this problem is well known and could be explained by the misalignment of mirrors. Would you know what problem it is? What would it be caused to? And what could we do to fix it?

Best regards,

Quentin and Mathieu

EVA Report – April 17th

EVA report 8

Date: 17/04/17

Weather: sunny, dry and hot, some clouds

Temperature: around 20°C

Location: around UTM NAD27 4251500-518500

Duration: 120 min

Team: Elke Mergny (Crew Geologist), Calogero Montedoro (Crew Biologist), Quentin Thomas (Crew Astronomer)

Purpose:

–          Geomatics study by 3D Mapping.

–          Aerial photos of an EVA

Outputs:

–          Mapping almost complete, but equipment issue at the end of the EVA, will not be salvageable with the equipment in the Hab.

–          Very nice aerial views were made.

Crew Photos – April 17th

Preparing field experiments

 

The Martian

 

Mmmmh turkey

 

Muon experiment

 

Riding to EVA

 

Planting rice

Journalist Report – April 17th

Journalist report Sol 7 – 170417

Mathieu « Mitch » Vander Donckt

Crew Journalist and Scientist

Crew 178 – UCL to Mars

 

A new day arises on Mars. The moral of the team is high, everybody being busy with their experiment or discovering new places and landscapes during Extra Vehicular Activity. We landed a week ago, and in a way a routine set in. Same schedule for the meals, at midday we mostly eat the bread that I bake – bread or brick, it depends on the courtesy of my crewmates – and in the evening with all our powder and dehydrated food we can let our creativity loose. Sometimes you even need more imagination to figure out what is in your plate, but it usually is tasty. We make one expedition every day, in general with a different team, location, or objective. We need to be coordinated: the team going outside the hab cannot split up for security reasons and our time is limited, so every second on the field matters. Living together, we get to know each other better. Boss is undoubtedly good at making unfunny jokes, Quentin aka “Dips”, our crew astronomer, is incline to fall asleep when we don’t give him an electrical shock, and we are all impressed by how good Tarzan is at crashing is drone.

But of course, every day has its part of discoveries and advances – this is why we are here. Now I’m working in the Science Dome with Tarzan on hydroponic culture of rice. We were respectively chemist and biologist before being astronaut, and our different ways of thinking about the same experiment is a good illustration of how the diversity of background is beneficial when composing an astronaut crew. The goal of this experiment is to assess the impact of urea on our plantations, and evaluate the impact of using our waste water to grow our food. One more step towards the food autonomy of Mars!

Interestingly, we watched “The Martian” a few days ago on our humble homemade cinema room and found some similarities between the movie and us. Let’s hope that we will not forget one of our crew members when we go home next week…

EVA Report – April 17th

EVA report 8

Date: 17/04/17

Weather: sunny, dry and hot, some clouds

Temperature: around 20°C

Location: around UTM NAD27 4251500-518500

Duration: 120 min

Team: Elke Mergny (Crew Geologist), Calogero Montedoro (Crew Biologist), Quentin Thomas (Crew Astronomer)

Purpose:

–          Geomatics study by 3D Mapping.

–          Aerial photos of an EVA

Outputs:

–          Mapping almost complete, but equipment issue at the end of the EVA, will not be salvageable with the equipment in the Hab.

–          Very nice aerial views were made.

Sol Summary – April 17th

Crew 178 Daily Summary Report 17APR2017

 

MDRS Daily Summary Report for sol 7

 

Summary Title: 7th and 8th EVA

 

Mission Status:  Good

 

Sol Activity Summary:

 

Hab

8.00:   Breakfast

13.00: Lunch

14.30: Preparation of the last telecommunication relays’s EVA

 

GreenHab

/

 

ScienceDome

Experiment reports redaction

Muons experiment: tests on the scintillator

 

EVA

9.00 UTM NAD27 12S 4255500-518500

X-ray diffraction study: collection of data

Positioning of sound beacons for mapping of dangerous area

Topography mapping

 

14.00: UTM NAD27 12S 4251500-518500

Topography mapping

Issue with the probe

 

Weather:

Sunny

 

Crew Physical Status

Everyone is fine

 

Look Ahead Plan:

EVA + Experiments

 

 

Reports to be file: Operation Report, Daily Summary Report, Journalist Report, HSO Report

 

Support Requested: /

EVA Report – April 17th

EVA report 7

Date: 17/04/17

Weather: sunny, dry and hot

Temperature: around 20°C

Location: UTM NAD27 12S 4255500-518500

Duration: 120 min

Team: Damien Mertens (crew engineer), Elke Mergny (crew geologist), Calogero Montedoro (crew biologist), Mathieu Vander Donckt (crew journalist)

Purpose:

–          Mapping of two areas with a drone.

–          Telecommunication: installation of marking device that use tools to cartography the sound beacons placed next to dangerous reliefs and hence, to construct a map of risk areas

Outputs:

–          Mapping of two area’s.

–          The installation of the marking device succeeded.

–          Some interesting soil samples were collected.

Astronomy Report – April 16th

Astronomy Report

Name: Quentin Thomas, Mathieu Vander Donckt      Crew: 178
date: 14 April 2017

Sky Conditions: clear

Wind Conditions: windy, we closed the lower part of the shutter because it was shivering.

Observation Start Time: 1am

Observation End Time: 2am

Summary: The first goal was to set up the telescope for the first time. We couldn’t make any photography due to the problems described below.

Objects Viewed: Arcturus, Jupiter

Problems Encountered: impossible to focus. For each viewed star, what should have been a single point was stretched out. Jupiter was fuzzy. We guess this is due to the alignment of the primary mirror. Can you send us the procedure to follow? Thank you.

Also, we didn’t dare to open the lower part of the shutter because of the wind. It was shivering. We let the upper part open to set up the telescope for next nights. After encountering the focusing problem, we closed everything.

 

Best regards,

Quentin “Dips” Thomas, Crew Astronomer,

Mathieu “Mitch” Vander Donckt, Crew Scientist and Journalist.

 

Crew Photos – April 16th

Just out of the station

 

Coach and hoover

 

Experimenting

 

Quad squad

 

The Station

 

EVA Team